# A Survey of Fuzzing Open-Source Operating Systems

### v2

@@ Abstract.
Vulnerabilities in open-source operating systems (OSs) pose substantial security risks to software systems, making their detection crucial. While fuzzing has been an effective vulnerability detection technique in various domains, OS fuzzing (OSF) faces unique challenges due to OS complexity and multi-layered interaction, and has not been comprehensively reviewed. Therefore, this work systematically surveys the state-of-the-art OSF techniques, categorizes them based on the general fuzzing process, and investigates challenges specific to kernel, file system, driver, and hypervisor fuzzing. Finally, future research directions for OSF are discussed. GitHub: https://github.com/pghk13/Survey-OSF.
@@ 1. Introduction
Operating systems (OSs) play a crucial role in the operation of modern computer systems. They are responsible for managing the hardware and software resources of a computer. They serve as the cornerstone for maximizing hardware resource utilization and ensuring software system stability. Specifically, open-source OSs, such as Linux, Android, FreeRTOS, Zephyr, etc., have become particularly favored in most fields like autonomous driving (Baidu, 2024; Chen et al., 2022b), robotics (ROS2, 2024), cloud services (Li et al., 2015), and Web of Things (Liu, 2023; Sun et al., 2023) for the advantages of transparency, customizability, and community-driven innovation, leading to their anticipated dominance in future computer systems.
However, as the scale of open-source OS codebases continues to expand, the security threats they pose have become increasingly alarming, raising significant concerns about the secure operation of software systems. According to National Vulnerability Database (NVD) (National Institute of Standards and Technology (2024), NIST), more than 9,300 vulnerabilities have been discovered in the kernels of mainstream open-source OSs (including Linux, Android, FreeBSD, OpenBSD and Zephyr) since 2004. In particular, the number of vulnerabilities surges to 3,300 in the single year of 2024, an increase that is 10 times higher than the previous year. Similarly, the Common Vulnerabilities and Exposures (CVE) (cve, 2024) has documented at least 7,600 vulnerabilities in the Linux kernel to date. Among them, NVD has disclosed at least 1,566 high-severity vulnerabilities and 156 critical vulnerabilities that require immediate remediation.
OS-level vulnerabilities are particularly concerning when compared to those in user-level applications. This is because such vulnerabilities can be further exploited, potentially allowing attackers to gain complete control over the OSs (ker, 2019; Xu et al., 2015; Zhang et al., 2015; Chen and Xing, 2019), leading to irreversible consequences. The staggering number of these vulnerabilities and the malicious outcomes they have caused have attracted significant attention from security researchers. As a result, there is a strong and growing interest in developing effective and efficient techniques to identify and mitigate these potential vulnerabilities, thereby aiding the continuous evolution of open-source OSs. One of the widely used techniques is fuzzing, which was first introduced by Miller et al. (Miller et al., 1990) in 1990, and has achieved notable success across various domains, e.g., compilers (Yang et al., 2023a), interpreters (Holler et al., 2012), and open-source software (Serebryany, 2017).
@@
However, to date, there has been no systematic review of OSF. While there are some survey papers on traditional software fuzzing (Eisele et al., 2022; Yun et al., 2022; Zhu et al., 2022; Manès et al., 2019; Li et al., 2018; Liang et al., 2018; Zhang et al., 2018; Mallissery and Wu, 2023; Böhme et al., 2020; Godefroid, 2020), they do not systematically introduce the general steps of OSF, and they also do not highlight the unique challenges that OSF faces compared to traditional software fuzzing. Therefore, to bridge the gap, a systematic review of the state-of-the-art OSF is essential, aiming to provide a comprehensive guideline on this topic.
@@
To systematically review OSF, we designed a comprehensive search process to identify existing high-quality research materials (see Section 2), including 58 state-of-the-art OSF papers and 4 open-source OSF tools. Through a thorough analysis of these materials, we uncovered the rising trend in OSF research and explained the reasons behind it. Then, we introduced the distinctive features of fuzzing techniques for the four OS layers (i.e., kernel, file system, driver, and hypervisor) from a high-level perspective, and summarized a general workflow of OSF, consisting of three core modules, i.e., input, fuzzing engine, and running environment (see Section 3). By analyzing the control flow and data flow between these modules, we further clarified the fundamental differences between OSF and traditional software fuzzing. Next, we conducted a comprehensive review of the state-of-the-art OSF with respect to seven key steps in the three core modules (see Section 4), aiming to systematically sort out the technological advancements in OSF and the associated complex domain knowledge. Meanwhile, we summarized the unique problems encountered by the four OS layers during fuzzing as well as the corresponding solutions (see Section 5). Finally, based on the findings of our review, we provided four future research directions in the OSF field (see Section 6).
@@ 2. Collection Strategy and Result
Following (Garousi and Mäntylä, 2016), we use a scientific and effective collection strategy (see Section 2.1) and present a detailed analysis of the collection result (see Section 2.2) to systematically review OSF.
@@ 2.1.1. Scientific Literature Review
Paper Filtering. We perform a manual assessment on the 56 candidate papers obtained from our database search to ensure their relevance and quality. Specifically, to determine whether each candidate paper is relevant to OSF and has high quality, we analyze the abstracts and introductions of these papers, following the inclusion and exclusion criteria formulated as follows.
  • Inclusion Criteria. IC1: papers that introduce the process of OSF; and IC2: papers that propose a technique of OSF.
  • Exclusion Criteria. EC1: survey papers or summary papers; EC2: papers that do not target fuzzing; EC3: papers that do not focus on fuzzing operating system and its components; and EC4: papers that have not been published in top-tier conferences or journals.
Specifically, for EC1, such survey papers are discussed in Section 3.1 for a comparison with our survey; and for EC4, we discuss these top-tier conferences and journals in Section 2.2. Through our manual assessment, we remove 13 papers, resulting in 43 papers.
Backward & Forward Snowballing. To reduce the risk of missing relevant papers, we perform both backward and forward snowballing (Wohlin, 2014) on the 43 papers. In backward snowballing, we check the references in these papers to obtain candidate papers, while in forward snowballing, we use Google Scholar to locate candidate papers that cite these papers. For these candidate papers obtained by snowballing, we also apply the same inclusion and exclusion criteria to identify relevant papers. Finally, we add 15 new relevant papers, resulting in a final set of 58 papers.
Full-Text Analysis. We download all the resulting 58 papers, and conduct a full-text analysis to identify the fuzzing target (i.e., the operating system and its components) and the proposed fuzzing technique. After reading all these papers, we classify them to form our survey (see Section 4 and 5).
@@ 2.2. Collection Result Analysis
Publication Venues. Figure 2(a) shows the distribution of all the papers across the publication venues. The 58 papers are published across 17 top-tier venues in four domains, i.e., security, software engineering, computer architecture, and computer storage systems. Specifically, (i) most of the papers, up to 79%, are published in security venues such as USENIX Security Symposium, ACM Conference on Computer and Communications Security (CCS), IEEE Symposium on Security and Privacy (S&P), and Network and Distributed System Security Symposium (NDSS); (ii) 12% of the papers are published in software engineering venues such as ACM Symposium on Operating Systems Principles (SOSP), International Symposium on Software Testing and Analysis (ISSTA), and International Conference on Software Engineering (ICSE); (iii) there are 4 papers, accounting for 7%, published in computer architecture venues, one each in European Conference on Computer Systems (EuroSys), USENIX Annual Technical Conference (ATC), ACM Transactions on Embedded Computing Systems (TECS), and IEEE Transactions on Computer-Aided Design of Integrated Circuits and Systems (TCAD); and (iv) since operating system is related to storage system, there is 1 paper published in computer storage system venues (i.e., ACM Transactions on Storage (TOS)). It can be conclude that the application of fuzzing techniques to operating systems spans multiple fields of computer science.
Publication Years and Target OS Layers. Figure 2(b) presents the number of papers published in each year as well as the distribution across the target OS layers in each year. Overall, the number of OSF papers shows a general ascending trend from 2015 to 2024. It is evident that interest in OSF has been escalating year by year, which indicates increased attention in leveraging fuzzing to uncover deeper security vulnerabilities in open-source OS. Notice that since the papers from 2024 have not been fully surveyed yet, Figure 2(b) only shows the number of OSF publications till August 2024. Moreover, most papers focus on the fuzzing of kernel. Particularly, Google’s fuzzing framework Syzkaller (Google, 2015) launched in 2015. It provides researchers with a foundational kernel fuzzing engine, and has made a substantial impact on kernel fuzzing. However, using kernel fuzzing interfaces (i.e., between user and kernel space) fails to uncover deep security vulnerabilities in other OS layers (e.g., verification chain checks in driver). Therefore, an increasing number of papers realized the necessity of fuzzing other OS layers (i.e., file systems, drivers, and hypervisors) since 2020.
@@ 3.1. Classification Dimension
Existing fuzzing surveys classify the literature by one of the four dimensions, i.e., 1) the amount of information the fuzzer requires or uses (black-, white-, and gray-box) (Manès et al., 2019; Li et al., 2018; Liang et al., 2018; Godefroid, 2020; Böhme et al., 2020), 2) the strategy employed for seed update (mutation- and generation-based) (Saavedra et al., 2019; Schumilo et al., 2020, 2021; Myung et al., 2022; Pan et al., 2021; Cesarano et al., 2023; Bulekov et al., 2022), 3) the research gaps of integrating advanced techniques into traditional fuzzing workflows (Wang et al., 2020; Zhang et al., 2018; Wang and Zhou, 2020; Zhu et al., 2022; Mallissery and Wu, 2023), and 4) the PUT (Yun et al., 2022; Eisele et al., 2022). Here, we adopt the fourth dimension as the guiding taxonomy, as the first three dimensions primarily focus on the general study of fuzzing techniques and lack the specificity required for analyzing OSF. In contrast, the PUT-based classification provides a comprehensive framework for systematically reviewing the general workflow of OSF while focusing on the unique challenges posed by the different OS layers (i.e., different PUTs).
@@ 3.2. OS Layers
Figure 3. The Interfaces of Each OS Layer ( 1, 2, 3 and 4 denote the fuzzing interfaces for Kernel, Driver, File System and Hypervisor respectively; and 1 is also the fuzzing interface for Driver and File System).
@@
Kernel. Kernel is one of the most critical systems in an OS because it manages essential resources such as processes and memory for the entire system and provides a unified programming interface for user-space programs to interact with hardware resources. All other system types rely on and operate on top of the kernel. Consequently, vulnerabilities in the kernel can be maliciously exploited, potentially causing significant damage to the systems running on it. As illustrated in Figure 3, to perform kernel fuzzing, a fuzzer triggers the kernel’s code paths by switching from user space to kernel space through syscall 1 (a.k.a. system call) (Koopman et al., 1997). Therefore, the core task of a kernel fuzzer is to generate various test cases (referred to as seeds) by combining syscalls provided by the kernel to continuously trigger the kernel’s code paths. For example, the darker sequences in Figure 3 denote branches covered by test cases, while the lighter circles denote code blocks yet to be covered.
@@
File System. As a core system service within an OS, file system is essential for tasks such as reading, writing, managing, and scheduling files, as well as ensuring data consistency during system crashes. Most file systems, like ext4 (Cao et al., 2007), XFS (and Inc(2018), SGI), Btrfs (Rodeh et al., 2013) and F2FS (Lee et al., 2015), operate within the OS kernel. Therefore, approaches used in kernel fuzzing can be adapted for file system fuzzing. However, approaches based purely on syscalls often yield numerous invalid results because the system state is predominantly influenced by metadata in file system operations. In contrast, operations on regular file data through syscalls like read() and write() contribute little to identifying file system vulnerabilities. Hence, an effective file system fuzzer typically combines sequences of file operation-related syscalls 1 with images 3 where metadata has been changed. As shown in Figure 3, the altered disk image is mounted to the file system’s partition using privileged commands (Schumilo et al., 2017; Kim et al., 2020b), becoming the new target for fuzzing. Kernel-provided syscalls are then used to conduct read, write, management, and scheduling operations, facilitating a comprehensive fuzzing of the file system.
@@
Driver. Drivers are responsible for communication and control between the user/kernel and hardware devices. They act as the bridge between hardware and OS, managing tasks such as device initialization, data transfer, and interrupt handling, ensuring that user-space applications can interact with underlying hardware (Chou et al., 2001; Palix et al., 2011; Stoep and Tolvanen, 2018). In driver fuzzing, the primary objective is to uncover vulnerabilities in device drivers across the initialization, data communication, and control stages. Since drivers can receive operation requests from both user space and hardware devices, they expose a broader attack surface compared to the kernel or other kernel subsystems (Beniamini, 2017a, b; Chang, 2017; Davis, 2011; Nohl, 2014). As illustrated in Figure 3, the two main attack surfaces for driver fuzzing are syscalls 1 and peripheral interfaces 2. Unlike kernel fuzzers, syscall-based driver fuzzers focus more on syscalls that directly operate on device files, e.g., read(), write(), seek(), ioctl(), etc.
@@
Hypervisor. In environments with highly heterogeneous hardware, OSs inherently lack the capability to manage and schedule heterogeneous computing resources, such as those used in industries like railways, avionics, and automotive systems (Cinque et al., 2022). To enable the concurrent execution of multiple OSs while maintaining secure isolation between them in such heterogeneous resource environments (RTCA, 1992; International Organization for Standardization, 2011), hypervisors, a.k.a. Virtual Machine Monitors (VMMs), leverage virtualization technique to partition hardware resources (e.g., CPU, memory, and disk space) into multiple virtual partitions (Popek and Goldberg, 1974; Cinque et al., 2022; Cilardo et al., 2022). When a guest OS, which runs within a Virtual Machine (VM), accesses or operates hardware devices, it triggers a VM-exit event that transfers these privileged operations to the hypervisor for hardware emulation. Therefore, the objective of hypervisor fuzzing is to accurately emulate the behavior of the guest OS when accessing and operating these virtualized hardware components, which serves as the primary entry point for implementing hypervisor fuzzing trials. As illustrated in Figure 3, the main interfaces involved include I/O channels and hypercalls 4.

### v3

@@ Abstract.
Vulnerabilities in open-source operating systems (OSs) pose substantial security risks to software systems, making their detection crucial. While fuzzing has been an effective vulnerability detection technique in various domains, OS fuzzing (OSF) faces unique challenges due to OS complexity and multi-layered interaction, and has not been comprehensively reviewed. Therefore, this work systematically surveys the state-of-the-art OSF techniques, categorizes them based on the general fuzzing process, and investigates challenges specific to kernel, file system, driver, and hypervisor fuzzing. Finally, future research directions for OSF are discussed.
@@ 1. Introduction
Operating systems (OSs) play a crucial role in the operation of modern computer systems. They are responsible for managing the hardware and software resources of a computer. They serve as the cornerstone for maximizing hardware resource utilization and ensuring software system stability. Specifically, open-source OSs, such as Linux, Android, FreeRTOS, Zephyr, etc., have become particularly favored in most fields like autonomous driving, robotics, cloud services, and Web of Things for the advantages of transparency, customizability, and community-driven innovation, leading to their anticipated dominance in future computer systems.
However, as the scale of open-source OS codebases continues to expand, the security threats they pose have become increasingly alarming, raising significant concerns about the secure operation of software systems. According to National Vulnerability Database (NVD) (NVD), more than 9,300 vulnerabilities have been discovered in the kernels of mainstream open-source OSs (including Linux, Android, FreeBSD, OpenBSD and Zephyr) since 2004. In particular, the number of vulnerabilities surges to 3,300 in the single year of 2024, an increase that is 10 times higher than the previous year. Similarly, the Common Vulnerabilities and Exposures (CVE) (cve) has documented at least 7,600 vulnerabilities in the Linux kernel to date. Among them, NVD has disclosed at least 1,566 high-severity vulnerabilities and 156 critical vulnerabilities that require immediate remediation.
OS-level vulnerabilities are particularly concerning when compared to those in user-level applications. This is because such vulnerabilities can be further exploited, potentially allowing attackers to gain complete control over the OSs (Xu2015FromCT; Zhang2015AndroidRA; chen2019slake), leading to irreversible consequences. The staggering number of these vulnerabilities and the malicious outcomes they have caused have attracted significant attention from security researchers. As a result, there is a strong and growing interest in developing effective and efficient techniques to identify and mitigate these potential vulnerabilities, thereby aiding the continuous evolution of open-source OSs. One of the widely used techniques is fuzzing, which was first introduced by Miller et al. (Miller1990AnES) in 1990, and has achieved notable success across various domains, e.g., compilers (Yang2023WhiteboxCF), interpreters (Holler2012FuzzingWC), and open-source software (Serebryany2017OSSFuzzG).
@@
However, to date, there has been no systematic review of OSF. While there are some survey papers on traditional software fuzzing (Eisele2022EmbeddedFA; Yun2022FuzzingOE; Zhu2022FuzzingAS; Mans2018TheAS; Li2018FuzzingAS; Liang2018FuzzingSO; Zhang2018SurveyOD; Mallissery2023DemystifyTF; bohme2020fuzzing; Godefroid2020), they do not systematically introduce the general steps of OSF, and they also do not highlight the unique challenges that OSF faces compared to traditional software fuzzing. Therefore, to bridge the gap, a systematic review of the state-of-the-art OSF is essential, aiming to provide a comprehensive guideline on this topic.
@@
To systematically review OSF, we designed a comprehensive search process to identify existing high-quality research materials (see Section 2), including 60 state-of-the-art OSF papers and 4 open-source OSF tools. Through a thorough analysis of these materials, we uncovered the rising trend in OSF research and explained the reasons behind it. Then, we introduced the distinctive features of fuzzing techniques for the four OS layers (i.e., kernel, file system, driver, and hypervisor) from a high-level perspective, and summarized a general workflow of OSF, consisting of three core modules, i.e., input, fuzzing engine, and running environment (see Section 3). Next, we conducted a comprehensive review of the state-of-the-art OSF with respect to seven key steps in the three core modules (see Section 4), aiming to systematically sort out the technological advancements in OSF and the associated complex domain knowledge. Meanwhile, we summarized the unique problems encountered by the four OS layers during fuzzing as well as the corresponding solutions (see Section LABEL:Section5). Finally, based on our findings, we provided four future research directions in the OSF field (see Section LABEL:Section7).
@@ 2. Collection Strategy and Result
Following (garousi2016systematic), we use a scientific and effective collection strategy (see Section 2.1) and present a detailed analysis of the collection result (see Section 2.2) to systematically review OSF.
@@ 2.1.1. Scientific Literature Review
Paper Filtering. We perform a manual assessment on the 56 candidate papers obtained from our database search to ensure their relevance and quality. Specifically, to determine whether each candidate paper is relevant to OSF and has high quality, we analyze the abstracts and introductions of these papers, following the inclusion and exclusion criteria formulated as follows. We use two inclusion criteria, i.e., IC1: papers that introduce the process of OSF, and IC2: papers that propose a technique of OSF. We use four exclusion criteria, i.e., EC1: survey papers or summary papers, EC2: papers that do not target fuzzing, EC3: papers that do not focus on fuzzing operating system and its components, and EC4: papers that have not been published in top-tier conferences or journals. Specifically, for EC1, such survey papers are discussed in Section 3.1 for a comparison with our survey; and for EC4, we discuss these top-tier conferences and journals in Section 2.2. Through our manual assessment, we remove 13 papers, resulting in 43 papers.
Backward & Forward Snowballing. To reduce the risk of missing relevant papers, we perform both backward and forward snowballing (wohlin2014guidelines) on the 43 papers. In backward snowballing, we check the references in these papers to obtain candidate papers, while in forward snowballing, we use Google Scholar to locate candidate papers that cite these papers. For these candidate papers obtained by snowballing, we also apply the same inclusion and exclusion criteria to identify relevant papers. Finally, we add 17 new relevant papers, resulting in a final set of 60 papers.
Full-Text Analysis. We download all the resulting 60 papers, and conduct a full-text analysis to identify the fuzzing target (i.e., the operating system and its components) and the proposed fuzzing technique. After reading all these papers, we classify them to form our survey (see Section 4 and LABEL:Section5).
@@ 2.2. Collection Result Analysis
Publication Venues. Figure 2(a) shows the distribution of all papers across the publication venues. The 60 papers are published across 18 top-tier venues in four domains, i.e., security, software engineering, computer architecture, and computer storage systems. Specifically, (i) most of the papers, up to 77%, are published in security venues such as USENIX Security Symposium (USEC), ACM Conference on Computer and Communications Security (CCS), IEEE Symposium on Security and Privacy (S&P), and Network and Distributed System Security Symposium (NDSS); (ii) 11% are published in software engineering venues such as ACM Symposium on Operating Systems Principles (SOSP), International Symposium on Software Testing and Analysis (ISSTA), and International Conference on Software Engineering (ICSE); (iii) there are 5 papers, accounting for 10%, published in computer architecture venues, one each in European Conference on Computer Systems (EuroSys), USENIX Annual Technical Conference (ATC), ACM Transactions on Embedded Computing Systems (TECS), IEEE Transactions on Computer-Aided Design of Integrated Circuits and Systems (TCAD), and International Conference on Architectural Support for Programming Languages and Operating Systems (ASPLOS); and (iv) since operating system is related to storage system, there is 1 paper published in computer storage system venues (i.e., ACM Transactions on Storage (TOS)). It can be conclude that application of fuzzing techniques to operating systems spans multiple fields of computer science.
Publication Years and Target OS Layers. Figure 2(b) presents the number of papers published in each year as well as the distribution across the target OS layers in each year. Overall, the number of OSF papers shows a general ascending trend from 2015 to 2024. It is evident that interest in OSF has been escalating year by year, which indicates increased attention in leveraging fuzzing to uncover deeper security vulnerabilities in open-source OS. Notice that since the papers from 2024 have not been fully surveyed yet, Figure 2(b) only shows the number of OSF publications till August 2024. Moreover, most papers focus on the fuzzing of kernel. Particularly, Google’s fuzzing framework Syzkaller (Syzkaller) launched in 2015. It provides researchers with a foundational kernel fuzzing engine, and has made a substantial impact on kernel fuzzing. However, using kernel fuzzing interfaces (i.e., between user and kernel space) fails to uncover deep security vulnerabilities in other OS layers (e.g., verification chain checks in driver). Therefore, an increasing number of papers realized the necessity of fuzzing other OS layers (i.e., file systems, drivers, and hypervisors) since 2020.
@@ 3.1. Classification Dimension
Existing fuzzing surveys classify the literature by one of the four dimensions, i.e., 1) the amount of information the fuzzer requires or uses (black-, white-, and gray-box) (Mans2018TheAS; Li2018FuzzingAS; Liang2018FuzzingSO; Godefroid2020; bohme2020fuzzing), 2) the strategy employed for seed update (mutation- and generation-based) (Saavedra2019ARO; schumilo2020hyper; schumilo2021nyx; myung2022mundofuzz; pan2021V-shuttle; cesarano2023iris; bulekov2022morphuzz), 3) the research gaps of integrating advanced techniques into traditional fuzzing workflows (Wang2019ASR; Zhang2018SurveyOD; Wang2020SoKTP; Zhu2022FuzzingAS; Mallissery2023DemystifyTF), and 4) the PUT (Yun2022FuzzingOE; Eisele2022EmbeddedFA). Here, we adopt the fourth dimension as the guiding taxonomy, as the first three dimensions primarily focus on the general study of fuzzing techniques and lack the specificity required for analyzing OSF. In contrast, the PUT-based classification provides a comprehensive framework for systematically reviewing the general workflow of OSF while focusing on the unique challenges posed by the different OS layers (i.e., different PUTs).
@@ 3.2. OS Layers
Figure 3. Interfaces of Each OS Layer (where 1, 2, 3 and 4 denote the fuzzing interfaces for Kernel, Driver, File System and Hypervisor respectively; and 1 is also the fuzzing interface for Driver and File System); and General Workflow of OSF (which has three modules, i.e., input, fuzzing engine, and running environment)
@@
Kernel. Kernel is one of the most critical systems in an OS because it manages essential resources such as processes and memory for the entire system and provides a unified programming interface for user-space programs to interact with hardware resources. All other system types rely on and operate on top of the kernel. Consequently, vulnerabilities in the kernel can be maliciously exploited, potentially causing significant damage to the systems running on it. As illustrated in Figure 3(a), to perform kernel fuzzing, a fuzzer triggers the kernel’s code paths by switching from user space to kernel space through syscall 1 (a.k.a. system call) (Koopman1997ComparingOS). Therefore, the core task of a kernel fuzzer is to generate various test cases (referred to as seeds) by combining syscalls provided by the kernel to continuously trigger the kernel’s code paths. For example, the darker sequences in Figure 3(a) denote branches covered by test cases, while the lighter circles denote code blocks yet to be covered.
@@
File System. As a core system service within an OS, file system is essential for tasks such as reading, writing, managing, and scheduling files, as well as ensuring data consistency during system crashes. Most file systems, like ext4 (Cao2007Ext4TN), XFS (XFS2018), Btrfs (Rodeh2013BTRFSTL) and F2FS (Lee2015F2FSAN), operate within the OS kernel. Therefore, approaches used in kernel fuzzing can be adapted for file system fuzzing. However, approaches based purely on syscalls often yield numerous invalid results because the system state is predominantly influenced by metadata in file system operations. In contrast, operations on regular file data through syscalls like read() and write() contribute little to identifying file system vulnerabilities. Hence, an effective file system fuzzer typically combines sequences of file operation-related syscalls 1 with images 3 where metadata has been changed. As shown in Figure 3(a), the altered disk image is mounted to the file system’s partition using privileged commands (schumilo2017kafl; Hydra2020finding), becoming the new target for fuzzing. Kernel-provided syscalls are then used to conduct read, write, management, and scheduling operations, facilitating a comprehensive fuzzing of the file system.
@@
Driver. Drivers are responsible for communication and control between the user/kernel and hardware devices. They act as the bridge between hardware and OS, managing tasks such as device initialization, data transfer, and interrupt handling, ensuring that user-space applications can interact with underlying hardware (Chou2001AnES; Palix2011FaultsIL; Stoep2018AndroidSecurity). In driver fuzzing, the primary objective is to uncover vulnerabilities in device drivers across the initialization, data communication, and control stages. Since drivers can receive operation requests from both user space and hardware devices, they expose a broader attack surface compared to the kernel or other kernel subsystems (Beniamini2017Part1; Beniamini2017Part2; Chang2017; Davis2011; NohlLell2014). As illustrated in Figure 3(a), the two main attack surfaces for driver fuzzing are syscalls 1 and peripheral interfaces 2. Unlike kernel fuzzers, syscall-based driver fuzzers focus more on syscalls that directly operate on device files, e.g., read(), write(), seek(), ioctl(), etc.
@@
Hypervisor. In environments with highly heterogeneous hardware, OSs inherently lack the capability to manage and schedule heterogeneous computing resources, such as those used in industries like railways, avionics, and automotive systems (Cinque2021VirtualizingMS). To enable the concurrent execution of multiple OSs while maintaining secure isolation between them in such heterogeneous resource environments (RTCA_DO_178C; ISO_26262_2011), hypervisors, a.k.a. Virtual Machine Monitors (VMMs), leverage virtualization technique to partition hardware resources (e.g., CPU, memory, and disk space) into multiple virtual partitions (Popek1974FormalRF; Cinque2021VirtualizingMS; Cilardo2021VirtualizationOM). When a guest OS, which runs within a Virtual Machine (VM), accesses or operates hardware devices, it triggers a VM-exit event that transfers these privileged operations to the hypervisor for hardware emulation. Therefore, the objective of hypervisor fuzzing is to accurately emulate the behavior of the guest OS when accessing and operating these virtualized hardware components, which serves as the primary entry point for implementing hypervisor fuzzing trials. As illustrated in Figure 3(a), the main interfaces involved include I/O channels and hypercalls 4.